OpenStudy (anonymous):

Hi how can I solve this trig identity? (sin(pi-x)cot(pi/2 - x) cos(2pi - x))/(tan(pi + x)tan(pi/2 + x)sin(-x)) = sinx ? can something like tan (a + b) be sin (a+ b)/cos(a + b)?

OpenStudy (anonymous):

yes, something like tan(a+b) can be sin(a+b)/cos(a+b)

OpenStudy (anonymous):

Make use of the odd function property of sin and the even function property of cos, these are: sin(-x) = -sin(x) and cos(-x)=cos(x) respectively. the fact that sin(x+Pi/2) = cos(x) and cos(x-Pi/2)=sin should help also

OpenStudy (anonymous):

Oh okay Thank you very much